# Cacophony: An Improved Contrastive Audio-Text Model

### v2

@@ VI-2 Dataset scale
We begin with the analysis of dataset scales in the first-stage training. Within each of the first three groups, (b-c), (d-e) and (f-g), the datasets used for the second-stage training are fixed, while the datasets used in the first-stage training are increased in size. Group (f-g) constitutes the extreme case: (f) uses the entire dataset for first-stage training and (g) bypasses the first stage entirely. Observations show an improvement in performance across all tasks and datasets as the first-stage training data expands. This evidences the positive impact of using larger datasets during the first-training stage on the model’s overall effectiveness. There is, however, an exception when comparing (a) with (b) and (c). There is no noticeable improvement in retrieval performance when comparing (a) and (c), and there is a decline in all evaluation metrics when additional data from ACAV2M is incorporated when comparing (a) and (b). This lack of improvement in (b) and (c) may be due to the evaluation dataset being out-of-distribution compared to the limited training dataset $V_{1}$ used in the second stage.
@@
We now analyze effects of data scales in the second-stage training, i.e., contrastive-captioning training. In the comparison of groups (c), (e), and (f), the audio encoder is initialized with weights from training on the $V_{3}$ dataset. During the second stage, these groups are trained on $V_{1}$, $V_{2}$, and $V_{3}$, respectively. When evaluating on AudioCaps, there is a significant increase in recall when integrating ACAV2M into the second stage training ($V_{2}$), but this improvement appears to plateau upon the inclusion of AudioSet ($V_{3}$). However, on Clotho, the retrieval performance consistently improves with increasing dataset scale, without observation of plateau, showing the effectiveness of increasing data scale.
@@
In experiments (a), (d), and (f), we increase dataset scale simultaneously for both training stages and observe a gradual enhancement in retrieval performance on the Clotho test dataset. However, a significant improvement in retrieval performance on the AudioCaps test dataset is observed only when integrating AudioSet (group (f)) into both training stages.
@@
We find scaling up the datasets in both training stages tends to lead to improved model performance. The pre-pretraining step with MAE contributes to the better retrieval performance for the audio-text model. Therefore, collecting more unlabeled audio or using larger versions of the ACAV dataset for pre-pretraining could potentially yield further gains in downstream performance.
@@
VI-3 Architectural design
In our ablation study focused on neural network architecture, we examine the effect of incorporating a captioning head, the use of SAM, and using a classifier pre-pretrained audio encoder on audio-text retrieval task presented in Table XI. For these experiments, we use the full-size dataset v3 in Tab X. We further demonstrate the adaptability of our audio encoder in handling varying audio lengths.
Captioning Objective. By comparing experiments (a) and (b) in Tab. XI, we observe a marginal improvement across both datasets when the captioning objective is included. We hypothesize that the captioning objective facilitates the learning of more fine-grained audio representations, which aligns with the findings of BLIP [26], CoCa [25] and CapCa [65].
Use of SAM. By comparing (a) and (c), we see that SAM significantly improves the accuracy on the audio-text retrieval task for both datasets. However, the benefits of SAM come at the expense of two sequential gradient computations during each training step; as mentioned earlier, there are methods that reduce SAM computational overhead that could be integrated into our training in future work.
Flexibility to Length. Lastly, we explore the capability of our model to process longer audio lengths at inference. We start by evaluating the audio-text retrieval task on the Clotho dataset, where audio sample durations range from 15 to 30 seconds. We compare the performance between group (a), which uses the entire length of the audio, and group (e), which randomly samples patches from the full audio to match the training duration. We find that using longer sequence lengths leads to an improvement in retrieval recall, even though our training involves only 10-second time-frequency patches. Then, we extend our investigation to include long sequence classification tasks from the HEAR benchmark, specifically focusing on the GTZAN [66] and VoxLingua [67] tasks, shown in Tab. XIII. Our findings reveal that leveraging the full length of the audio samples consistently yields higher classification accuracy compared to using fixed length-patch subsets, which is consistent with findings in FLIP [22].
@@ VII Conclusions
In this paper, we propose both the curation of a large paired audio-text dataset and modeling improvements to enhance the training of joint audio-language models. By automatically captioning unlabeled audio data and refining noisy existing captions with language models, we curate a dataset of 3.9 million text-audio pairs. Our modeling approach consists of a two-stage training strategy: The first stage involves training an audio encoder with an MAE objective, while the second stage combines contrastive learning with an auxiliary captioning task. Our final model, Cacophony, achieves state-of-the-art performance on audio-text retrieval tasks and competitive performance on other downstream tasks.
To evaluate the impacts of our dataset curation and modeling strategies, we conduct a series of ablation studies. Notably, we observe a positive influence of our modeling techniques and a positive role of dataset scaling in both training phases. The improvements seen from scaling MAE pre-pretraining data are particularly relevant, as MAE training leverages audio-only data, facilitating further dataset scaling in future work.
Acknowledgments
We are immensely grateful to the Google TPU Research Cloud (TRC) for generously providing the computational resources vital to our project. Their support has been invaluable.
We thank the Freesound team from Pompeu Fabra University for providing us with the scraping API. We thank the University of Rochester Goergen Institute for Data Science (GIDS) seed funding program. We thank LAION CLAP team for collecting open source datesets and for generously sharing them with the research community.

### v3

@@ VI-B Dataset scale
We first study the effect of dataset scale in the first-stage training. In Table IX), each of the three groups, (b-c), (d-e) and (f-g), uses a different dataset in the first-stage training but the same dataset in the second-stage training. In particular, Group (f-g) constitutes the extreme case: (f) uses the entire dataset in the first-stage training and (g) bypasses the first stage entirely. Observations show a performance improvement across all tasks and datasets as the first-stage training data expands. This evidences the positive impact of using larger datasets during the first-training stage on the model’s overall effectiveness. There are, however, two exceptions: There is no noticeable improvement in retrieval performance when comparing (a) and (c), and there is a decline in all evaluation metrics when additional data from ACAV2M is incorporated when comparing (a) and (b). This lack of improvement in (b) and (c) compared to (a) may be due to the evaluation dataset being out-of-distribution of the limited training dataset $V_{1}$ used in the second stage.
@@
We now analyze effects of data scale in the second-stage training, i.e., contrastive-captioning training. In the comparison of experiments (c), (e), and (f), the audio encoder is initialized with weights from training on the $V_{3}$ dataset. During the second stage, these groups are trained on $V_{1}$, $V_{2}$, and $V_{3}$, respectively. When evaluating on AudioCaps, there is a significant increase from (c) to (e) in recall when integrating ACAV2M ($V_{2}$) into the second-stage training, but this improvement appears to plateau from (e) to (f) upon the inclusion of AudioSet ($V_{3}$). However, on Clotho, the retrieval performance consistently improves with increasing dataset scale from (c) to (e) and (f), without observation of a plateau, showing the effectiveness of increasing data scale.
@@
In experiments (a), (d), and (f), we increase the dataset scale simultaneously for both training stages and observe a gradual improvement on retrieval performance on the Clotho test dataset. However, a significant improvement in retrieval performance on AudioCaps is observed only when integrating AudioSet (group (f)) into both training stages.
@@
We find scaling up the datasets in both stages tends to lead to improved model performance. The MAE training step contributes to the better retrieval performance for the audio-text model. Therefore, collecting more unlabeled audio or using larger versions of the ACAV dataset for audio encoder MAE training could potentially yield further gains in downstream performance.
@@
VI-C Architectural design
Regarding the neural network architecture, we examine the effect of incorporating a captioning head, SAM optimization, and fixed length processing window on audio-text retrieval tasks (see Table XI). For these experiments, we use the full-size dataset $V_{3}$ in both training stages, i.e., setting (f) in Table IX. We further demonstrate the adaptability of our audio encoder in handling varying audio lengths.
VI-C1 Captioning Objective.
By comparing experiments (a) and (b) in Table XI, we observe a marginal improvement across both test sets when the captioning objective is included. We hypothesize that the captioning objective facilitates the learning of more fine-grained audio representations, which aligns with the findings of BLIP [27], CoCa [26] and CapCa [62].
VI-C2 Use of SAM.
By comparing (a) and (c), we see that SAM significantly improves the accuracy on the audio-text retrieval task for both datasets. However, the benefits of SAM come at the expense of two sequential gradient computations during each training step; as mentioned earlier, there are methods for reducing SAM computational overhead that could be integrated in future work.
VI-C3 Flexibility to Length.
Lastly, we explore the capability of our model to process longer audio during inference. We start by evaluating the audio-text retrieval task on the Clotho dataset, where audio sample duration ranges from 15 to 30 seconds. We compare the performance between (a), which uses the entire length of the audio, and (e), which randomly samples 10-second spectrogram MAE patches from the full audio to match the training duration. We find that using longer sequence lengths leads to an improvement in retrieval recall, even though our training only uses 10-second time-frequency patches. Then, we extend our investigation to include long sequence classification tasks from the HEAR benchmark, specifically focusing on the GTZAN [63] and VoxLingua [64] tasks, as shown in Table XII. Our findings reveal that leveraging the full length of the audio samples consistently yields higher classification accuracy compared to using fixed length patches, which is consistent with findings in FLIP [23].
@@ VII Conclusions
In this paper, we presented dataset and model improvements for contrastive audio-language models. By automatically captioning unlabeled audio data and refining noisy existing captions with language models, we curated a dataset of 3.9 million text-audio pairs. Our modeling approach consists of a two-stage training strategy: The first stage involves training an audio encoder with an MAE objective, while the second stage combines contrastive learning with an auxiliary captioning task. Our final model, Cacophony, achieves state-of-the-art performance on audio-text retrieval tasks and competitive performance on other downstream tasks.
We conducted a series of ablation studies to evaluate the effects of our dataset curation and modeling strategies. Notably, we observed positive effects of our modeling techniques and dataset scaling in both training phases. The improvement seen from scaling MAE pretraining data is particularly relevant, as MAE training leverages audio-only data and can easily scale further in future work.
While our work demonstrates significant advancements in audio-text modeling, it is important to acknowledge several limitations. We observe weaker performance in fine-grained classification and captioning tasks, suggesting the need for more detailed, higher-quality captions in our training data. In particular, the reliance on a single-modality audio captioner for synthesizing a large portion of our data may limit the diversity and richness of our dataset. Integrating multimodal captioning synthesis techniques, such as those involving the visual modality, could potentially address this limitation. Furthermore, our model’s relatively weak performance on speech-related tasks indicates a need for more comprehensive speech data integration. Finally, the persistent modality gap between audio and text embeddings, for our proposed model and other audio-text models, remains a challenge, potentially limiting performance in tasks requiring tight cross-modal alignment.